# Can LLM-based Financial Investing Strategies Outperform the Market in Long Run?

### v3

@@ Appendix A Data Collection
The financial news dataset, initially compiled by Dong et al. (2024), comprises 15.7 million records pertaining to 4,775 S&P 500 companies, spanning the years 1999 to 2023. We have organised the news by aligning it with the respective companies and indexing it by date.
@@ B.1. Timing-based Strategies
This category includes FinMem (Yu et al., 2023) and FinRobot (Yang et al., 2024). We acknowledge other works, such as FinCon (Yu et al., 2024) and MarketSenseAI (Fatouros et al., 2025), but they are not (yet) open-source, which prevents us from generating backtesting results.
@@
We implement and cover several well-known traditional rule-based (indicator-based) investing strategies, such as Buy and Hold, Simple Moving Average Crossover, Weighted Moving Average Crossover, ATR Band, Bollinger Bands (Bollinger, 2002), Trend Following (Wilcox et al., 2005), and Turn of the Month (McConnell and Xu, 2008). These strategies typically rely on one or multiple technical indicators or domain-based rules to generate timely buy/sell signals, aiming to exploit identifiable market patterns or anomalies.
@@
In contrast to fixed rules or indicator-based triggers, these strategies rely on data-driven models—often statistical or neural network forecasters—to predict future price movements. Specifically, they buy or hold if an uptrend is indicated and sell (or go short) otherwise. This can be viewed as a relatively naive application of ML/DL forecasters, but it is widely used as a benchmark method for such models. Although one could consider the forecast output as a type of “indicator”, the reliance on predictive algorithms capable of uncovering complex patterns sets these methods apart from purely rule-based approaches. We include the well-known ARIMA (Box and Jenkins, 1990) and XGBoost (Chen and Guestrin, 2016) in this category and also cover forecasters based on LLMs, differentiating them from the LLM investors.
@@
We also implement widely used RL algorithms for financial markets, including Advantage Actor-Critic (A2C), Proximal Policy Optimisation (PPO), Twin Delayed Deep Deterministic Policy Gradient (TD3), and Soft Actor-Critic (SAC), utilising the FinRL framework (Liu et al., 2022, 2024). Each agent learns investing policies by interacting with a simulated trading environment based on the OpenAI Gym API, using real historical market data.
@@ B.2. Selection-based Strategies
Following the well-documented momentum factor (Muller and and, 2010), this strategy selects the stocks with the highest recent price appreciation. For each candidate stock, we calculate a momentum score based on its historical price data. Specifically, the score is the percentage return over a “momentum period” (e.g., 100 trading days), but we exclude the most recent “skip period” (e.g., 21 trading days) from the calculation. This practice is common in momentum strategies to avoid the “short-term reversal” effect (Carhart, 1997). The score for a given stock is calculated as shown in Equation 1. $t$ is the selection date. All candidate stocks are then ranked in descending order by this score, and the top-$k$ stocks (e.g., $k=5$) are selected.
@@
This strategy is based on the “volatility effect” anomaly, where low-volatility stocks have been empirically shown to generate higher risk-adjusted returns (Blitz and Vliet, 2007). For each candidate stock, we measure its historical volatility over a recent “look-back period” (e.g., 21 trading days). The volatility is calculated as the standard deviation of its weekly log returns within this period. We use weekly returns ($\ln(P_{t}/P_{t-5})$) rather than daily returns to smooth out daily noise. Candidate stocks are then ranked in ascending order by their calculated volatility, and the top-$k$ stocks with the lowest volatility are selected for the portfolio.
@@
Unlike the single-factor methods above, the FinCon Selection Agent (Yu et al., 2024) aims to construct a diversified portfolio by explicitly considering both performance and inter-stock correlation. Its selection process is more sophisticated:
@@ Appendix D Extra Results on Selective Symbols
Tables 2 and 7 further substantiate our findings by highlighting the performance instability of FinMem and FinAgent when extending evaluation periods even marginally. Specifically, extending the evaluation by just two months beyond the originally reported periods (Yu et al., 2023) results in notable inconsistencies in critical performance metrics. It should be noted that the results for the LLM strategies are retrieved from Yu et al. (2024), while the traditional rule-based results presented are based on our implementations.
@@
Table 7. Backtest performance of traditional rule-based (indicator-based) strategies and FinCon over the selective period (2022-10-05 to 2023-06-10), as presented in Yu et al. (2024), evaluated using four metrics: cumulative return (CR), Sharpe ratio (SPR), annual volatility (AV), and maximum drawdown (MDD). The best metrics are highlighted in red, while the second best are marked in blue. “-” metrics across the board indicate no trade signals were triggered.
@@ Appendix E Technical Details
The backtesting framework and traditional rule-based strategies in FINSABER are implemented using BackTrader [^9] https://www.backtrader.com/ and Papers With Backtest [^10] https://paperswithbacktest.com/. Reinforcement learning-based methods are implemented using FinRL (Liu et al., 2021). FINSABER supports two operational modes: “LLM” mode and “BT” mode. The “LLM” mode is tailored for strategies that leverage multi-modal inputs, including financial news and regulatory filings. In contrast, the “BT” mode is built directly on BackTrader, offering robust support for traditional rule-based strategies while maintaining a familiar interface to facilitate easy migration from standard BackTrader workflows.
@@ Appendix F Comparative Drawdown Analysis via Underwater Plots
The third row, depicting the 2008 Global Financial Crisis, provides the most critical insight into the agents’ flaws. While a single stock is expected to be more volatile than the index during a crash, the LLM strategies, particularly FinMem, catastrophically amplify this downside risk. For DE, the FinMem strategy’s drawdown approaches -75%, a far more severe loss than the SPX benchmark’s -50%. Rather than providing any form of risk mitigation, the agents appear to make pro-cyclical decisions that accelerate losses. The FinAgent strategy, true to its more conservative nature, often mitigates some of these losses relative to FinMem, yet it still fails to generate a positive outcome. For instance, while its drawdown for CNX is shallower than FinMem’s, it remains severe and prolonged. This relative outperformance is insufficient and aligns with our market regime analysis (§7), which finds that both agents are poorly calibrated for bear markets and ultimately succumb to losses (Cont, 2001).

### v4

@@ Appendix A Data Collection
The financial news dataset, initially compiled by 10.1145/3637528.3671629, comprises 15.7 million records pertaining to 4,775 S&P 500 companies, spanning the years 1999 to 2023. We have organised the news by aligning it with the respective companies and indexing it by date.
@@ B.1. Timing-based Strategies
This category includes FinMem (yu2023finmemperformanceenhancedllmtrading) and FinRobot (yang2024finrobotopensourceaiagent). We acknowledge other works, such as FinCon (yu2024fincon) and MarketSenseAI (fatouros2025marketsenseai20enhancingstock), but they are not (yet) open-source, which prevents us from generating backtesting results.
@@
We implement and cover several well-known traditional rule-based (indicator-based) investing strategies, such as Buy and Hold, Simple Moving Average Crossover, Weighted Moving Average Crossover, ATR Band, Bollinger Bands (bollinger2002bollinger), Trend Following (Wilcox2009trendfollowing), and Turn of the Month (turnofthemonth2008). These strategies typically rely on one or multiple technical indicators or domain-based rules to generate timely buy/sell signals, aiming to exploit identifiable market patterns or anomalies.
@@
In contrast to fixed rules or indicator-based triggers, these strategies rely on data-driven models—often statistical or neural network forecasters—to predict future price movements. Specifically, they buy or hold if an uptrend is indicated and sell (or go short) otherwise. This can be viewed as a relatively naive application of ML/DL forecasters, but it is widely used as a benchmark method for such models. Although one could consider the forecast output as a type of “indicator”, the reliance on predictive algorithms capable of uncovering complex patterns sets these methods apart from purely rule-based approaches. We include the well-known ARIMA (10.5555/574978) and XGBoost (10.1145/2939672.2939785) in this category and also cover forecasters based on LLMs, differentiating them from the LLM investors.
@@
We also implement widely used RL algorithms for financial markets, including Advantage Actor-Critic (A2C), Proximal Policy Optimisation (PPO), Twin Delayed Deep Deterministic Policy Gradient (TD3), and Soft Actor-Critic (SAC), utilising the FinRL framework (finrl-2022; finrl-2024). Each agent learns investing policies by interacting with a simulated trading environment based on the OpenAI Gym API, using real historical market data.
@@ B.2. Selection-based Strategies
Following the well-documented momentum factor (Muller31032010), this strategy selects the stocks with the highest recent price appreciation. For each candidate stock, we calculate a momentum score based on its historical price data. Specifically, the score is the percentage return over a “momentum period” (e.g., 100 trading days), but we exclude the most recent “skip period” (e.g., 21 trading days) from the calculation. This practice is common in momentum strategies to avoid the “short-term reversal” effect (on-persistence-1997). The score for a given stock is calculated as shown in Equation 1. $t$ is the selection date. All candidate stocks are then ranked in descending order by this score, and the top-$k$ stocks (e.g., $k=5$) are selected.
@@
This strategy is based on the “volatility effect” anomaly, where low-volatility stocks have been empirically shown to generate higher risk-adjusted returns (volatility-effect). For each candidate stock, we measure its historical volatility over a recent “look-back period” (e.g., 21 trading days). The volatility is calculated as the standard deviation of its weekly log returns within this period. We use weekly returns ($\ln(P_{t}/P_{t-5})$) rather than daily returns to smooth out daily noise. Candidate stocks are then ranked in ascending order by their calculated volatility, and the top-$k$ stocks with the lowest volatility are selected for the portfolio.
@@
Unlike the single-factor methods above, the FinCon Selection Agent (yu2024fincon) aims to construct a diversified portfolio by explicitly considering both performance and inter-stock correlation. Its selection process is more sophisticated:
@@ Appendix D Extra Results on Selective Symbols
Tables 2 and 7 further substantiate our findings by highlighting the performance instability of FinMem and FinAgent when extending evaluation periods even marginally. Specifically, extending the evaluation by just two months beyond the originally reported periods (yu2023finmemperformanceenhancedllmtrading) results in notable inconsistencies in critical performance metrics. It should be noted that the results for the LLM strategies are retrieved from yu2024fincon, while the traditional rule-based results presented are based on our implementations.
@@
Table 7. Backtest performance of traditional rule-based (indicator-based) strategies and FinCon over the selective period (2022-10-05 to 2023-06-10), as presented in yu2024fincon, evaluated using four metrics: cumulative return (CR), Sharpe ratio (SPR), annual volatility (AV), and maximum drawdown (MDD). The best metrics are highlighted in red, while the second best are marked in blue. “-” metrics across the board indicate no trade signals were triggered.
@@ Appendix E Technical Details
The backtesting framework and traditional rule-based strategies in FINSABER are implemented using BackTrader [^9] https://www.backtrader.com/ and Papers With Backtest [^10] https://paperswithbacktest.com/. Reinforcement learning-based methods are implemented using FinRL (liu2021finrl). FINSABER supports two operational modes: “LLM” mode and “BT” mode. The “LLM” mode is tailored for strategies that leverage multi-modal inputs, including financial news and regulatory filings. In contrast, the “BT” mode is built directly on BackTrader, offering robust support for traditional rule-based strategies while maintaining a familiar interface to facilitate easy migration from standard BackTrader workflows.
@@ Appendix F Comparative Drawdown Analysis via Underwater Plots
The third row, depicting the 2008 Global Financial Crisis, provides the most critical insight into the agents’ flaws. While a single stock is expected to be more volatile than the index during a crash, the LLM strategies, particularly FinMem, catastrophically amplify this downside risk. For DE, the FinMem strategy’s drawdown approaches -75%, a far more severe loss than the SPX benchmark’s -50%. Rather than providing any form of risk mitigation, the agents appear to make pro-cyclical decisions that accelerate losses. The FinAgent strategy, true to its more conservative nature, often mitigates some of these losses relative to FinMem, yet it still fails to generate a positive outcome. For instance, while its drawdown for CNX is shallower than FinMem’s, it remains severe and prolonged. This relative outperformance is insufficient and aligns with our market regime analysis (§7), which finds that both agents are poorly calibrated for bear markets and ultimately succumb to losses (cont2001empirical).